a vaccine refrigerator is set to 5 degrees . but temparature may vary upto 3 degrees. which absolute value equation can be used to find the maximam and minimum temperatures of the refrigerator ?
a) 1x1 –5 = 3 b) 1x-51 =3 c) 1x1 –3 =5 d) 1x-31 =5

Answer by stanbon(75887)

a vaccine refrigerator is set to 5 degrees . but temparature may vary upto 3 degrees. which absolute value equation can be used to find the maximam and minimum temperatures of the refrigerator ?
a) 1x1 –5 = 3 b) 1x-51 =3 c) 1x1 –3 =5 d) 1x-31 =5
------------------
The distance from 5 degrees is 3
|x-5|=3
==========================================
Cheers,
Stan H.